# Identifying Knowledge Editing Types in Large Language Models

### v2

@@
Abstract
@@
Knowledge editing has emerged as an efficient technology for updating the knowledge of large language models (LLMs), attracting increasing attention in recent years. However, there is a lack of effective measures to prevent the malicious misuse of this technology, which could lead to harmful edits in LLMs. These malicious modifications could cause LLMs to generate toxic content, misleading users into inappropriate actions. In front of this risk, we introduce a new task, Knowledge Editing Type Identification (KETI), aimed at identifying different types of edits in LLMs, thereby providing timely alerts to users when encountering illicit edits. As part of this task, we propose KETIBench, which includes five types of harmful edits covering most popular toxic types, as well as one benign factual edit. We develop four classical classification models and three BERT-based models as baseline identifiers for both open-source and closed-source LLMs. Our experimental results, across 42 trials involving two models and three knowledge editing methods, demonstrate that all seven baseline identifiers achieve decent identification performance, highlighting the feasibility of identifying malicious edits in LLMs. Additional analyses reveal that the performance of the identifiers is independent of the reliability of the knowledge editing methods and exhibits cross-domain generalization, enabling the identification of edits from unknown sources. All data and code are available in https://github.com/xpq-tech/KETI. Warning: This paper contains examples of toxic text.
1 Introduction
@@
Knowledge editing is an emerging technology designed to efficiently rectify errors or outdated knowledge in large language models (LLMs) (Yao et al., 2023). In recent years, it has garnered increasing attention (Wang et al., 2023). The primary aim of knowledge editing is to provide approaches for efficiently updating the knowledge within LLMs, thereby avoiding the costly process of retraining LLMs. However, there is a deficiency in effective measures to prevent the misuse of these techniques (Chen et al., 2024; Li et al., 2024b). Some malicious individuals may exploit knowledge editing to inject harmful information into LLMs, which does not align with human values. Users lacking discernment can be easily influenced by those problematic information, potentially leading them to engage in actions that contravene ethical standards and human values.
@@
Current research begins preliminary exploration into detecting edited knowledge within LLMs (Youssef et al., 2024). This study finds that simple classifiers can effectively distinguish between edited and non-edited knowledge. However, it fails to distinguish between various forms of edited knowledge, which is vital as it aids in discerning the intentions of editors and thus enables the effective alerting of users to potentially harmful and erroneous information that can arise from illicit edits of LLMs. To fill this gap, we propose a new task called Knowledge Editing Type Identification (KETI), which aims to identify the knowledge edited in LLMs and its corresponding type. Providing more fine-grained information about the knowledge editing allows users to differentiate between malicious and benign updates more effectively, enabling timely alerts to users when encountering illicit edits. The KETI task is examplified in Figure 1.
Figure 1: Illustration of the KETI task.
@@
In this study, we explore how to identify different types of edited knowledge. First, we propose a new benchmark for KETI called KETIBench. We construct a dataset containing both malicious and benign updates. We follow existing research on the safety of LLMs Zou et al. (2023); Wang et al. (2024a); Chen et al. (2024), considering misinformation injection, offensiveness injection, behavioral misleading injection, and bias injection as malicious updates, while benign updates are fact updating. Second, we design seven baseline identifiers for both open-source and closed-source LLMs. For open-source LLMs, we use the hidden states of the LLMs as features and employ four classical classification models to learn the patterns of hidden states for different types of edits. For closed-source LLMs, we use the query, output, and log probabilities of the output token as features and employ three BERT-based models to learn the input-output features of different types of edits. Third, we select three knowledge editing methods—FT-M (Zhang et al., 2024), GRACE (Hartvigsen et al., 2023), and UnKE (Deng et al., 2024)—and conduct experiments on two LLMs, Llama3.1-7B-Instruct (Dubey et al., 2024) and Llama2-13B-Chat (Touvron et al., 2023), under both open-source and closed-source settings. Experimental results demonstrate that seven baseline identifiers achieve decent performance on KETIBench, with an average F1 score of 0.789 across 42 sets of experiments. Despite this, the performance of the identifiers still requires improvement, as a 20% error identification rate still poses a significant risk. We also provide some useful insights on how to better perform the KETI task. 1) We find a weak correlation between performance of identifiers and the reliability of knowledge editing; 2) We demonstrate the ability of baseline identifiers to identify edits from unknown sources through cross-domain experiments; 3) We identify the key factors determining the performance of the identifiers through in-depth analysis and ablation studies. We find that the richness of feature information determines the performance of the identifier. In open-source LLMs, the hidden states in the later layers that carry richer semantic information are more helpful for the identifier to identify different types of edits. In closed-source LLMs, more information from the output tokens contributes to the identifier’s performance. Our research provides a preliminary exploration into identifying knowledge editing types in LLMs.
2 Related Work
@@
Knowledge editing aims to provide efficient approaches to update the knowledge within LLMs, a rapidly emerging area that has gained increasing attention (Yao et al., 2023; Zhang et al., 2024; Wang et al., 2023). In recent years, various knowledge editing methods (Meng et al., 2022b; Li et al., 2023; Mitchell et al., 2022; Hartvigsen et al., 2023; Deng et al., 2024; Wang et al., 2024b; Li et al., 2024a) and benchmarks (Wang et al., 2024c; Rosati et al., 2024; Zhong et al., 2023) have been proposed. However, the ethical and moral risks associated with knowledge editing should also be acknowledged. Researchers have found that backdoors and harmful information can be implanted into LLMs through knowledge editing (Li et al., 2024b; Chen et al., 2024). Therefore, there is an urgent need for measures to mitigate the risks posed by malicious edits. To this end, we propose the KETI task, which aims to identify the knowledge editing types. Results from various baseline identifiers on KETIBench show that different types of knowledge editing in both open-source and closed-source LLMs can be identified. Our preliminary exploration makes it possible to effectively mitigate the risks posed by the misuse of knowledge editing. Notably, the most relevant work to ours is DEED (Youssef et al., 2024); however, DEED only considers the distinction between edits and non-edits, without further exploring different types of edits.
3 Knowledge Editing
@@
4 Identifying Knowledge Editing Types in Large Language Models
4.1 KETI Task
@@ 4.2 Data Construction
The preconditions for KETI is that an LLM has already been edited with a series of knowledge, primarily harmful information, along with a small amount of beneficial knowledge (i.e., fact updating). Inspired by previous works (Zou et al., 2023; Wang et al., 2024a; Chen et al., 2024), we consider five types of knowledge in KETIBench: Fact updating, Misinformation injection, Offensiveness injection, Behavioral misleading injection, and Bias injection.
Figure 2: Data Statistics. Turquoise indicates samples used for editing, while other colors represent samples not used for editing. Different colors denote subcategories within the samples not used for editing.
@@
We first collect Behavioral misleading, Offensiveness, Misinformation, and Bias data from Advbench (Zou et al., 2023). Advbench contains 576 harmful strings, but these strings do not have corresponding queries, rephrased queries, or specific categories. Therefore, we use GPT-4o to generate these details for the strings. We show the prompts in Appendix A.2. Second, we enrich the number of Misinformation and Bias samples using Misinformation Injection and Bias Injection from the Editing Attack dataset (Chen et al., 2024). Third, we select samples from the zsRE dataset (Mitchell et al., 2022) where the answer length exceeds 50 characters for use in Fact updating. The rationale for selecting samples with an answer length greater than 50 is that the average character length of the object $o$ in other types of knowledge is near 50, and we aim to minimize the impact of varying object $o$ lengths on the experimental conclusions. Finally, to simulate real-world scenarios, we consider to include non-edited knowledge in KETIBench. To this end, we sample a proportional number of non-edited samples from Safeedit (Wang et al., 2024a) and zsRE, matching the six types. Finally, we obtain 1,432 samples, with dataset statistics shown in the figure 2. We also show some data samples in Appendix A.2. We then divide KETIBench into a test set and a training set according to a 3:7 ratio based on different editing types.
4.3 Baseline Identifiers
@@
Current LLMs can be categorized into two types based on whether their weights are open-source: open-source LLMs and closed-source LLMs. The information we can access from these two types of LLMs differs. For open-source LLMs, we can access all information, including weights, hidden states generated during inference, and output data. In contrast, for closed-source LLMs, only the output data is accessible. Therefore, we describe different baseline identifiers for each type of LLM.
4.3.1 Baseline Identifiers for open-source LLMs
@@
Since the information accessible in closed-source LLMs is a subset of that in open-source LLMs, the identifiers for close-source LLMs are also applicable to open-source LLMs. Here, we only consider weights and hidden states, leaving output data to be considered under closed-source LLMs. When the original model weights are available, it is easy to determine whether the weights have been modified by comparing them, and thus determining whether the model has been edited is trivial. However, model weights are static features, and relying solely on this static feature does not provide insight into the internal state of LLMs with respect to the knowledge prompts $\mathcal{P}$, making it difficult to identify the type of knowledge edits. In contrast, hidden states, as dynamic features, aggregate information from both the knowledge prompts and the weights (Geva et al., 2023), offering richer information. Therefore, we design identifiers using the hidden states generated by the knowledge prompts within LLMs as features. Furthermore, the hidden states $h\in R^{1\times h_{\text{dim}}}$ of the last token in the last layer of autoregressive LLMs aggregate dense information (Geva et al., 2023; Youssef et al., 2024), so we use only these hidden states as features. In the remainder of this paper, we use “hidden states” to refer to the hidden states of the last token in the last layer.
@@
  • Linear Discriminant Analysis (LDA) (Hastie et al., 2009): It is a supervised learning algorithm designed to achieve dimensionality reduction and classification by projecting data into a lower-dimensional space, maximizing between-class variance while minimizing within-class variance. For each class $c$, the discriminant function of LDA is:
    $\delta_{c}(h)=h^{T}\Sigma^{-1}\mathbf{\mu}_{c}-\frac{1}{2}\mathbf{\mu}_{c}^{T}\Sigma^{-1}\mathbf{\mu}_{c}+\log\mathbb{P}(y=c)$ (1)
@@
    $\mathbb{P}(y=c|h)=\frac{\exp(h^{T}\mathbf{w}_{c}+b_{c})}{1+\sum_{c^{\prime}=1}^{|c|}\exp(h^{T}\mathbf{w}_{c^{\prime}}+b_{c^{\prime}})}$ (2)
@@
  • Linear Classifier (Linear): It learns the weight $\mathbf{W}\in R^{hidden\_dim\times c}$ and bias $b$ used for classification during training:
    $y=h\mathbf{W}+b$ (3)
@@
    $y=\left(\sigma\left(\sigma\left(h\mathbf{W_{1}}+b_{1}\right)\mathbf{W_{2}}+b_{2}\right)\mathbf{W_{3}}+b_{3}\right)$ (4)
    where $\sigma$ is the activation function; $\mathbf{W_{1}}\in R^{hidden\_dim\times m_{1}}$, $\mathbf{W_{1}}\in R^{m_{1}\times m_{2}}$, and $\mathbf{W_{1}}\in R^{m_{2}\times c}$ are weights; $b_{1}$, $b_{2}$, and $b_{3}$ are bias.
4.3.2 Baselines identifiers for closed-source LLMs
The most direct information produced by closed-source LLMs after inference is the output text $t$. Additionally, they may also output the top-K log probabilities $l=[l_{1},...,l_{n}]\in R^{n\times K}$ of n tokens. For example, the OpenAI API  [^1] https://platform.openai.com/docs/api-reference/chat/create can return top-20 log probabilities of output tokens. We can therefore use these information as features. To better identify different editing types, we can also use the knowledge prompts $\mathcal{P}=[p_{1},\dots]$ as features. Considering that the primary feature of closed-source LLMs is text, we mainly build baseline identifiers based on BERT (Devlin, 2018).

### v3

@@
Abstract.
Warning: This paper contains examples of toxic text.
@@ Abstract.
Knowledge editing has emerged as an efficient technique for updating the knowledge of large language models (LLMs), attracting increasing attention in recent years. However, there is a lack of effective measures to prevent the malicious misuse of this technique, which could lead to harmful edits in LLMs. These malicious modifications could cause LLMs to generate toxic content, misleading users into inappropriate actions. In front of this risk, we introduce a new task, Knowledge Editing Type Identification (KETI), aimed at identifying different types of edits in LLMs, thereby providing timely alerts to users when encountering illicit edits. As part of this task, we propose KETIBench, which includes five types of harmful edits covering the most popular toxic types, as well as one benign factual edit. We develop five classical classification models and three BERT-based models as baseline identifiers for both open-source and closed-source LLMs. Our experimental results, across 92 trials involving four models and three knowledge editing methods, demonstrate that all eight baseline identifiers achieve decent identification performance, highlighting the feasibility of identifying malicious edits in LLMs. Additional analyses reveal that the performance of the identifiers is independent of the reliability of the knowledge editing methods and exhibits cross-domain generalization, enabling the identification of edits from unknown sources. All data and code are available in https://github.com/xpq-tech/KETI.
Knowledge Editing, Model Editing, Large Language Model, Security of Knowledge Editing
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737001isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Natural language generationccs: Security and privacy Malware and its mitigation
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15481859.
1. INTRODUCTION
@@
Knowledge editing is an emerging technique designed to efficiently rectify errors or outdated knowledge in large language models (LLMs) (Yao et al., 2023; Zhang et al., 2024a; Wang et al., 2023). In recent years, it has garnered increasing attention (Wang et al., 2023). The primary aim of knowledge editing is to provide approaches for efficiently updating the knowledge within LLMs, thereby avoiding the costly process of retraining LLMs. However, there is a deficiency in effective measures to prevent the misuse of these techniques (Chen et al., 2024; Li et al., 2024a). Some malicious individuals may exploit knowledge editing to inject harmful information into LLMs, which does not align with human values. Users lacking discernment can be easily influenced by those problematic information, potentially leading them to engage in actions that contravene human values.
Figure 1. Illustration of the KETI task. After an LLM has undergone both benign and harmful edits, it becomes difficult for users to distinguish whether the content generated by the LLM is a result of harmful edits. However, KETI can use technical methods to distinguish whether the edits are harmful.
[…]
Illustration of the KETI task.
@@
Current research begins preliminary exploration into detecting edited knowledge within LLMs (Youssef et al., 2024). This study finds that simple classifiers can effectively distinguish between edited and non-edited knowledge. However, merely detecting whether knowledge has been edited offers limited value for downstream decision-making. Users struggle to assess the trustworthiness of edited outputs, while developers cannot infer whether the edits reflect benign intentions or malicious manipulation. To fill this gap, we introduce a new task called Knowledge Editing Type Identification (KETI), which aims to identify the knowledge editing types in LLMs. Providing more fine-grained information about the knowledge editing allows users to differentiate between malicious and benign updates more effectively, enabling timely alerts to users when encountering illicit edits. The KETI task is examplified in Figure 1.
@@
In this study, we explore how to identify different types of edited knowledge. First, we propose a new benchmark for KETI called KETIBench. We construct a dataset containing both malicious and benign updates. We follow existing research on the safety of LLMs (Zou et al., 2023; Wang et al., 2024c; Chen et al., 2024), considering misinformation injection, offensiveness injection, behavioral misleading injection, and bias injection as malicious updates, while benign updates are fact updating. Second, we design eight baseline identifiers for both open-source and closed-source LLMs. For open-source LLMs, we use the hidden states of the LLMs as features and employ five classical classification models to learn the patterns of hidden states for different types of edits. For closed-source LLMs, we use the query, output, and log probabilities of the output token as features and employ three BERT-based models to learn the input-output features of different types of edits. Third, we select three knowledge editing methods—FT-M (Zhang et al., 2024a), GRACE (Hartvigsen et al., 2023), and UnKE (Deng et al., 2024)—and conduct experiments on four LLMs, Llama3.1-7B-Instruct (Dubey et al., 2024), Llama2-13B-Chat (Touvron et al., 2023), Qwen2.5-7B-Instruct (Team, 2024) and GPT2-XL under both open-source and closed-source settings. Experimental results demonstrate that seven baseline identifiers achieve decent performance on KETIBench, with an average F1 score of 0.745 across 92 sets of experiments. Despite this, the performance of the identifiers still requires improvement, as a 25% error identification rate still poses a significant risk. We also provide some useful insights on how to better perform the KETI task. 1) We find a weak correlation between the performance of identifiers and the reliability of knowledge editing; 2) We demonstrate the ability of baseline identifiers to identify edits from unknown sources through cross-domain experiments; 3) We identify the drawbacks of baseline identifiers through in-depth analyses, providing fine-grained insights into how to better design identifiers: a) The baseline identifiers introduced in this paper fail to fully leverage both the hidden state features of open-source LLMs and the output features of closed-source LLMs to distinguish between edited and non-edited knowledge. b) The hidden state features of open-source LLMs are complex and intertwined, making baseline identifiers challenging to interpret. c) The strongest identifier for closed-source LLMs has limitations in both semantic understanding and probabilistic feature interpretation. We provide a preliminary exploration into identifying knowledge editing types in LLMs.
In summary, our contributions are as follows:
  • Introduction of the KETI task and KETIBench: We introduce KETIBench, a benchmark that includes a dataset containing both malicious and benign edits. Malicious edits are categorized into misinformation, offensiveness, behavioral misleading, and bias injections, while benign edits involve fact updating.
  • Introduction and evaluation of baseline identifiers: We introduce eight baseline identifiers for open-source and closed-source LLMs. We evaluate them on four LLMs, edited using three knowledge editing methods — FT-M, GRACE, and UnKE. The results show that the identifiers perform decently, but there is still room for further improvement.
  • Key insights: 1) There is a weak correlation between the performance of the identifiers and the reliability of knowledge editing; 2) Baseline identifiers generalize well to edits from unknown sources, as shown in cross-domain experiments; 3) Baseline identifiers face limitations in leveraging hidden state features of open-source LLMs and output features of closed-source LLMs, struggle with the complexity of open-source model features, and lack robust semantic understanding and probabilistic feature interpretation for closed-source models.
2. RELATED WORK
@@
Knowledge editing aims to provide efficient approaches to update the knowledge within LLMs, a rapidly emerging area that has gained increasing attention (Yao et al., 2023; Zhang et al., 2024a; Wang et al., 2023). In recent years, various knowledge editing methods (Meng et al., 2022b; Li et al., 2023; Mitchell et al., 2022; Hartvigsen et al., 2023; Deng et al., 2024; Wang et al., 2024a; Li et al., 2024b; Fang et al., 2024; Gu et al., 2024; Ma et al., 2024) and benchmarks (Wang et al., 2024b; Rosati et al., 2024; Zhong et al., 2023; Huang et al., 2024; Zhang et al., 2024b) have been proposed. However, the ethical and moral risks associated with knowledge editing should also be acknowledged. Researchers have found that backdoors and harmful information can be implanted into LLMs through knowledge editing (Li et al., 2024a; Chen et al., 2024). Therefore, there is an urgent need for measures to mitigate the risks posed by malicious edits. To this end, we propose the KETI task, which aims to identify the knowledge editing types. Results from various baseline identifiers on KETIBench show that different types of knowledge editing in both open-source and closed-source LLMs can be identified. Our preliminary exploration makes it possible to effectively mitigate the risks posed by the misuse of knowledge editing. Notably, the most relevant work to ours is DEED (Youssef et al., 2024); however, DEED only considers the distinction between edits and non-edits, without further exploring different types of edits.
3. KNOWLEDGE EDITING
@@
4. KETIBENCH & BASELINES
4.1. KETI Task
@@ 4.2. Data Construction
The preconditions for KETI is that an LLM has already been edited with a series of knowledge, primarily harmful edits, along with a small amount of beneficial knowledge (i.e., fact updating). Inspired by previous works (Zou et al., 2023; Wang et al., 2024c; Chen et al., 2024), we consider five types of knowledge in KETIBench: Fact updating, Misinformation injection, Offensiveness injection, Behavioral misleading injection, and Bias injection.
@@
We first collect Behavioral misleading, Offensiveness, Misinformation, and Bias data from Advbench (Zou et al., 2023). Advbench contains 576 harmful strings, but these strings do not have corresponding queries, rephrased queries, or specific categories. Therefore, we use GPT-4o to generate these details for the strings. We show the prompts in Appendix C. Second, we enrich the number of Misinformation and Bias samples using Misinformation Injection and Bias Injection from the Editing Attack dataset (Chen et al., 2024). Third, we select samples from the zsRE dataset (Mitchell et al., 2022) where the answer length exceeds 50 characters for use in Fact updating. The rationale for selecting samples with an answer length greater than 50 is that the average character length of the object $o$ in other types of knowledge is near 50, and we aim to minimize the impact of varying object $o$ lengths on the experimental conclusions. Finally, to simulate real-world scenarios, we consider to include non-edited knowledge in KETIBench. To this end, we sample a proportional number of non-edited samples from Safeedit (Wang et al., 2024c) and zsRE, matching the six types. Finally, we obtain 1,432 samples, with dataset statistics shown in Table 1. We also show some data samples in Appendix D. We then divide KETIBench into a test set and a training set according to a 3:7 ratio based on different editing types.
Behav. Bias Offen. Misin. Fact Total
Edit 307 178 144 307 221 1157
Non-Edited 79 45 37 76 38 275
Table 1. Data statics. Behav. denotes Behavioral misleading; Offen. denotes Offensiveness; Misin. denotes Misinformation.
4.3. Baseline Identifiers
@@
Current LLMs can be categorized into two types based on whether their weights are open-source: open-source LLMs and closed-source LLMs. The information we can access from these two types of LLMs differs. For open-source LLMs, we can access all information, including weights, hidden states generated during inference, and output data. In contrast, for closed-source LLMs, only the output data is accessible. Therefore, we describe different baseline identifiers for each type of LLM. Notably, the open-source and closed-source scenarios both essentially involve identifying features of the same model. Therefore, the results of the BERT-based method for closed-source LLMs are equivalent to those for open-source LLMs.
4.3.1. Baseline Identifiers for open-source LLMs
@@
Since the information accessible in closed-source LLMs is a subset of that in open-source LLMs, the identifiers for closed-source LLMs are also applicable to open-source LLMs. Here, we only consider weights and hidden states, leaving output data to be considered under closed-source LLMs. When the original model weights are available, it is easy to determine whether the weights have been modified by comparing them, and thus determining whether the model has been edited is trivial. However, model weights are static features, and relying solely on this static feature does not provide insight into the internal state of LLMs with respect to the knowledge prompts $\mathcal{P}$, making it difficult to identify the type of knowledge edits. In contrast, hidden states, as dynamic features, aggregate information from both the knowledge prompts and the weights (Geva et al., 2023), offering richer information. Therefore, we design identifiers using the hidden states generated by the knowledge prompts within LLMs as features. Furthermore, the hidden states $h\in R^{1\times h_{\text{dim}}}$ of the last token in the last layer of autoregressive LLMs aggregate dense information (Geva et al., 2023; Youssef et al., 2024), so we use only these hidden states as features. In the remainder of this paper, we use “hidden states” to refer to the hidden states of the last token in the last layer.
@@
  • Linear Discriminant Analysis (LDA) (Hastie et al., 2009b): It is a supervised learning algorithm designed to achieve dimensionality reduction and classification by projecting data into a lower-dimensional space, maximizing between-class variance while minimizing within-class variance. For each class $c$, the discriminant function of LDA is:
    (1) $\delta_{c}(h)=h^{T}\Sigma^{-1}\mathbf{\mu}_{c}-\frac{1}{2}\mathbf{\mu}_{c}^{T}\Sigma^{-1}\mathbf{\mu}_{c}+\log\mathbb{P}(y=c)$
@@
  • AdaBoost (Adaptive Boosting) (Hastie et al., 2009a): It is previously introduced in DEED (Youssef et al., 2024). It is an ensemble learning method that combines multiple weak classifiers to form a strong classifier by iteratively focusing on the samples that are misclassified. At each iteration $t$, a weak classifier $h_{t}(h)$ is trained and assigned a weight $\alpha_{t}$ based on its performance:
    (2) $\alpha_{t}=\ln\left(\frac{1-\epsilon_{t}}{\epsilon_{t}}\right)$
    where $\epsilon_{t}$ is the weighted error rate of $h_{t}(h)$, defined as:
    (3) $\epsilon_{t}=\frac{\sum_{i=1}^{n}w_{i}^{(t)}\mathbb{I}(h_{t}(h_{i})\neq y_{i})}{\sum_{i=1}^{n}w_{i}^{(t)}}$
    Here, $w_{i}^{(t)}$ represents the weight of sample $i$ at iteration $t$, and $\mathbb{I}$ is the indicator function. The weights of the samples are updated to emphasize the misclassified ones:
    (4) $w_{i}^{(t+1)}=w_{i}^{(t)}\exp\left(\alpha_{t}\mathbb{I}(h_{t}(h_{i})\neq y_{i})\right)$
    The final prediction is made by taking a weighted majority vote of the weak classifiers:
    (5) $\hat{y}=\arg\max_{c}\sum_{t=1}^{T}\alpha_{t}\mathbb{I}(h_{t}(h)=c)$
    where $T$ is the total number of iterations, and $h_{t}(h)$ represents the prediction of the $t$-th weak classifier.
@@
    (6) $\mathbb{P}(y=c|h)=\frac{\exp(h^{T}\mathbf{w}_{c}+b_{c})}{1+\sum_{c^{\prime}=1}^{|c|}\exp(h^{T}\mathbf{w}_{c^{\prime}}+b_{c^{\prime}})}$
@@
  • Linear Classifier (Linear): It learns the weight $\mathbf{W}\in R^{h_{dim}\times c}$ and bias $b$ used for classification during training:
    (7) $y=h\mathbf{W}+b$
@@
    (8) $y=\left(\sigma\left(\sigma\left(h\mathbf{W_{1}}+b_{1}\right)\mathbf{W_{2}}+b_{2}\right)\mathbf{W_{3}}+b_{3}\right)$
    where $\sigma$ is the activation function; $\mathbf{W_{1}}\in R^{h_{dim}\times m_{1}}$, $\mathbf{W_{2}}\in R^{m_{1}\times m_{2}}$, and $\mathbf{W_{3}}\in R^{m_{2}\times c}$ are weight matrices, with hyperparameters controlling the intermediate sizes $m_{1}$ and $m_{2}$; $b_{1}$, $b_{2}$, and $b_{3}$ are bias.
Editor Source Identifier Llama3.1-8B-Instruct Llama2-13B-chat Qwen2.5-7B-Instruct GPT2-XL
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
FT-M Open LDA 0.712 0.714 0.711 0.790 0.789 0.789 0.694 0.701 0.694 0.661 0.674 0.659
AdaBoost 0.623 0.623 0.623 0.624 0.626 0.623 0.643 0.653 0.646 0.549 0.552 0.550
LogR 0.843 0.843 0.842 0.842 0.841 0.840 0.769 0.768 0.767 0.811 0.821 0.815
Linear 0.856 0.837 0.837 0.797 0.795 0.783 0.786 0.687 0.688 0.826 0.825 0.825
MLP 0.823 0.828 0.821 0.787 0.763 0.764 0.714 0.717 0.711 0.808 0.828 0.814
Close BERT-text only 0.808 0.756 0.771 0.772 0.735 0.747 0.828 0.822 0.821 0.814 0.822 0.817
BERT+SFLP 0.849 0.847 0.847 0.842 0.795 0.807 0.834 0.811 0.817 0.848 0.777 0.791
BERT+LSTM 0.863 0.861 0.861 0.815 0.825 0.818 0.821 0.836 0.809 0.841 0.841 0.837
GRACE Open LDA 0.778 0.778 0.778 0.775 0.780 0.777 0.748 0.731 0.733 0.616 0.632 0.620
AdaBoost 0.622 0.618 0.615 0.547 0.546 0.545 0.511 0.502 0.505 0.498 0.499 0.497
LogR 0.833 0.837 0.834 0.815 0.820 0.816 0.828 0.824 0.825 0.779 0.778 0.778
Linear 0.827 0.834 0.823 0.808 0.801 0.804 0.758 0.670 0.691 0.810 0.775 0.778
MLP 0.808 0.814 0.796 0.675 0.636 0.614 0.560 0.452 0.414 0.775 0.770 0.761
Close BERT-text only 0.772 0.767 0.763 0.797 0.781 0.781 0.809 0.794 0.792 0.797 0.780 0.779
BERT+SFLP 0.781 0.764 0.770 0.805 0.755 0.757 0.811 0.804 0.801 0.849 0.840 0.842
BERT+LSTM 0.770 0.765 0.762 0.813 0.813 0.801 0.793 0.794 0.790 0.801 0.811 0.802
UnKE Open LDA 0.783 0.783 0.782 0.741 0.750 0.744 0.763 0.764 0.763 0.692 0.705 0.696
AdaBoost 0.635 0.621 0.626 0.535 0.530 0.529 0.590 0.581 0.583 0.585 0.558 0.568
LogR 0.846 0.850 0.847 0.819 0.820 0.819 0.836 0.827 0.830 0.807 0.813 0.809
Linear 0.859 0.855 0.854 0.789 0.746 0.754 0.827 0.653 0.620 0.770 0.750 0.735
MLP 0.810 0.819 0.801 0.706 0.710 0.694 0.465 0.521 0.460 0.725 0.712 0.697
Close BERT-text only 0.797 0.802 0.796 0.798 0.769 0.778 0.734 0.754 0.737 0.742 0.723 0.721
BERT+SFLP 0.805 0.772 0.782 0.786 0.773 0.763 0.768 0.776 0.770 0.790 0.763 0.774
BERT+LSTM 0.831 0.818 0.818 0.798 0.798 0.797 0.762 0.764 0.756 0.808 0.782 0.782
Table 2. Results of baseline identifiers. Red background indicates the best performance for Open LLMs. Purple background indicates the best performance for Close LLMs. Bold indicates the best performance across all types of LLMs.
4.3.2. Baselines identifiers for closed-source LLMs
API Log Probabilities
OpenAI Top-20
DeepSeek Top-20
Claude Unsupported
Gemini Top-K
Grok Top-8
Table 3. Top-K log probability support across popular closed-source APIs. “Top-K” means the API supports access to Top-K log probabilities, though the exact value of K is unknown. “Unsupported” indicates that this feature is not available.
The most direct information produced by closed-source LLMs after inference is the output text $t$. Additionally, they may also output the top-K log probabilities $l=[l_{1},...,l_{n}]\in R^{n\times K}$ of n tokens. In practice, many closed-source models (i.e., APIs) expose Top-K logits. For instance, OpenAI and Gemini provide this functionality. We summarize our findings in Table 3. For example, the OpenAI API can return top-20 log probabilities of output tokens. We can therefore use these information as features. To better identify different editing types, we can also use the knowledge prompts $\mathcal{P}=[p_{1},\dots]$ as features. Considering that the primary feature of closed-source LLMs is text, we mainly build baseline identifiers based on BERT (Devlin, 2018).